OpenStudy (anonymous):

What is the equation of a circle with a center at (4, 0) and a radius of 1?

OpenStudy (anonymous):

(x-4)^2 +y^2=1 x^2 -8x +y^2 +16 -1=0 x^2 -8x +y^2 +15=0

OpenStudy (anonymous):

thanks @matricked :)

OpenStudy (anonymous):

welcome